# Market Making without Regret

### v1

@@ 3.2 $T^{\nicefrac{{2}}{{3}}}$ lower bound (IID+Lip+IV)
We recall that if $(\mathcal{X},\mathcal{F})$ is a measurable space and $X$ is a $\mathcal{X}$-valued random variable, we denote by $\mathbb{P}_{X}$ the push-forward probability measure of $\mathbb{P}$ induced by $X$ on $\mathcal{X}$, i.e., $\mathbb{P}_{X}[E]\coloneqq\mathbb{P}[X\in E]$, for any $E\in\mathcal{F}$. Also, with $\left\lVert\cdot\right\rVert_{\text{TV}}$ we denote the total variation norm for measures, and for any two probability measures $\mathbb{Q},\mathbb{Q}^{\prime}$ defined on the same sample space, we denote their Kullback-Leibler divergence by $\mathcal{D}_{\text{KL}}(\mathbb{Q},\mathbb{Q}^{\prime})$. Now, for each $k\in[K]$, using Pinsker’s inequality [65, Lemma 2.5] that upper bounds the difference in the total variation $\left\lVert\cdot\right\rVert_{\text{TV}}$ of two probability measures using their Kullback-Leibler divergence $\mathcal{D}_{\text{KL}}$, we have that
@@
$\mathcal{D}_{\text{KL}}\left(\mathbb{P}_{(M_{t+1},W_{t+1})\mid(M_{1},W_{1},\dots,M_{t},W_{t})=(m_{1},w_{1},\dots,m_{t},w_{t}))},\mathbb{P}_{(M_{t+1},W_{t+1}^{k})\mid(M_{1},W_{1}^{k},\dots,M_{t},W_{t}^{k})=(m_{1},w_{1},\dots,m_{t},w_{t}))}\right)\\ =\mathcal{D}_{\text{KL}}\left(\mathbb{P}_{(\mathbb{I}\{V_{t+1}\leq b\},\mathbb{I}\{a<V_{t+1}\})},\mathbb{P}_{(\mathbb{I}\{V_{t+1}^{k}\leq b\},\mathbb{I}\{a<V_{t+1}^{k}\})}\right)\cdot\bigg{(}\mathbb{I}\{(a,b)\in R^{\textbf{left}}_{k}\}+\sum_{j=1}^{k-1}\mathbb{I}\{(a,b)\in R^{\textbf{square}}_{k,j}\}\\ +\mathbb{I}\{(a,b)\in R^{\textbf{triangle}}_{k}\}+\sum_{i=1}^{k-1}\mathbb{I}\{(a,b)\in R^{\textbf{square}}_{i,k}\}+\mathbb{I}\{(a,b)\in R^{\textbf{top}}_{k}\}+\mathbb{I}\{(a,b)\in R^{\textbf{exploit}}_{k}\}\bigg{)}$
A direct verification (see Lemma C.3 in Appendix C) shows that, letting $c_{2}\coloneqq\nicefrac{{65}}{{9}}$, for each $k\in[K]$, and $i,j\in[K-1]$, if $(a,b)\in R^{\textbf{left}}_{k}\cup R^{\textbf{square}}_{k,j}\cup R^{\textbf{triangle}}_{k}\cup R^{\textbf{square}}_{i,k}\cup R^{\textbf{top}}_{k}$, then
@@
likewise (see, again, Lemma C.3 in Appendix C), letting $c_{1}\coloneqq\nicefrac{{2}}{{81}}$ such that, for each $k\in[K]$, if $(a,b)\in R^{\textbf{exploit}}_{k}$, then
@@
and notice that, for each $t\in[T-1]$ such that $t\geq 2$, using the chain rule for the Kullback-Leibler divergence [31, Theorem 2.5.3], we have that
@@
$\qquad+\int_{\left([0,1]\times\{0,1\}^{2}\right)^{t-1}}\mathcal{D}_{\text{KL}}\left(\mathbb{P}_{t,m_{1:t-1},w_{1:t-1}},\mathbb{P}^{k}_{t,m_{1:t-1},w_{1:t-1}}\right)\,\mathrm{d}\mathbb{P}_{(M_{1},W_{1},\dots,M_{t-1},W_{t-1})}(m_{1},w_{1},\dots,m_{t-1},w_{t-1})$

### v2

@@ Appendix B Additional Related Works
We recall that if $(\mathcal{X},\mathcal{F})$ is a measurable space and $X$ is a $\mathcal{X}$-valued random variable, we denote by $\mathbb{P}_{X}$ the push-forward probability measure of $\mathbb{P}$ induced by $X$ on $\mathcal{X}$, i.e., $\mathbb{P}_{X}[E]\coloneqq\mathbb{P}[X\in E]$, for any $E\in\mathcal{F}$. Also, with $\left\lVert\cdot\right\rVert_{\text{TV}}$ we denote the total variation norm for measures, and for any two probability measures $\mathbb{Q},\mathbb{Q}^{\prime}$ defined on the same sample space, we denote their Kullback-Leibler divergence by $\mathcal{D}_{\text{KL}}(\mathbb{Q},\mathbb{Q}^{\prime})$. Now, for each $k\in[K]$, using Pinsker’s inequality (Tsybakov, 2008, Lemma 2.5) that upper bounds the difference in the total variation $\left\lVert\cdot\right\rVert_{\text{TV}}$ of two probability measures using their Kullback-Leibler divergence $\mathcal{D}_{\text{KL}}$, we have that
@@
$\mathcal{D}_{\text{KL}}\big{(}\mathbb{P}_{(M_{t+1},W_{t+1})\mid(M_{1},W_{1},\dots,M_{t},W_{t})=(m_{1},w_{1},\dots,m_{t},w_{t}))},$
$\qquad\qquad\mathbb{P}_{(M_{t+1},W_{t+1}^{k})\mid(M_{1},W_{1}^{k},\dots,M_{t},W_{t}^{k})=(m_{1},w_{1},\dots,m_{t},w_{t}))}\big{)}$
$=\mathcal{D}_{\text{KL}}\left(\mathbb{P}_{(\mathbb{I}\{V_{t+1}\leq b\},\mathbb{I}\{a<V_{t+1}\})},\mathbb{P}_{(\mathbb{I}\{V_{t+1}^{k}\leq b\},\mathbb{I}\{a<V_{t+1}^{k}\})}\right)$
$\qquad\cdot\bigg{(}\mathbb{I}\{(a,b)\in R^{\textbf{left}}_{k}\}+\sum_{j=1}^{k-1}\mathbb{I}\{(a,b)\in R^{\textbf{square}}_{k,j}\}+\mathbb{I}\{(a,b)\in R^{\textbf{triangle}}_{k}\}$
$\qquad\qquad+\sum_{i=1}^{k-1}\mathbb{I}\{(a,b)\in R^{\textbf{square}}_{i,k}\}+\mathbb{I}\{(a,b)\in R^{\textbf{top}}_{k}\}+\mathbb{I}\{(a,b)\in R^{\textbf{exploit}}_{k}\}\bigg{)}$
A direct verification (see Lemma D.6) shows that, letting $c_{2}\coloneqq\nicefrac{{65}}{{9}}$, for each $k\in[K]$, and $i,j\in[K-1]$, if $(a,b)\in R^{\textbf{left}}_{k}\cup R^{\textbf{square}}_{k,j}\cup R^{\textbf{triangle}}_{k}\cup R^{\textbf{square}}_{i,k}\cup R^{\textbf{top}}_{k}$, then
@@
likewise (see, again, Lemma D.6), letting $c_{1}\coloneqq\nicefrac{{2}}{{81}}$ such that, for each $k\in[K]$, if $(a,b)\in R^{\textbf{exploit}}_{k}$, then
@@
and notice that, for each $t\in[T-1]$ such that $t\geq 2$, using the chain rule for the Kullback-Leibler divergence (Cover and Thomas, 2006, Theorem 2.5.3), we have that
@@
$\qquad+\int_{\left([0,1]\times\{0,1\}^{2}\right)^{t-1}}\mathcal{D}_{\text{KL}}\big{(}\mathbb{P}_{t,m_{1:t-1},w_{1:t-1}},\mathbb{P}^{k}_{t,m_{1:t-1},w_{1:t-1}}\big{)}$
$\qquad\qquad\qquad\qquad\,\mathrm{d}\mathbb{P}_{(M_{1},W_{1},\dots,M_{t-1},W_{t-1})}(m_{1},w_{1},\dots,m_{t-1},w_{t-1})$